A steel cylinder of cylinder of 500mm outside diameter and 200mm inside diameter is set in rotation
about its axis. If the cylinder is 900mm long, of density 7800kg/m3
, calculate the torque required to
give an angular acceleration of 0.5rad/s2
.
1
Expert's answer
2020-10-15T10:47:18-0400

Solution

Inner redius(b)=200 mm=0.2 m

Outer redius(a)=500 mm=0.5 m

Height of cylender(h)=900 mm=0.9 m

Dencity of cylender("\rho" )=7800 kg/m^3

Moment of inertia can be

"I=\pi h\rho \frac{(a^4-b^4)}{2}"

So by putting the value

"I=3.14\times 7800\times0.9\times \frac{(0.5^4-0.2^4)}{2}"

"I=671.2 kg m^2"

Torque can be given as

"\tau=I\alpha"

"\tau=671.2\times0.5=335.6 Nm"